### v2

# Dialectal Coverage And Generalization in Arabic Speech Recognition

Abstract
Developing robust automatic speech recognition (ASR) systems for Arabic, a language characterized by its rich dialectal diversity and often considered a low-resource language in speech technology, demands effective strategies to manage its complexity. This study explores three critical factors influencing ASR performance: the role of dialectal coverage in pre-training, the effectiveness of dialect-specific fine-tuning compared to a multi-dialectal approach, and the ability to generalize to unseen dialects. Through extensive experiments across different dialect combinations, our findings offer key insights towards advancing the development of ASR systems [^1] The code with pre-trained & finetuned model weights is available on GitHub: https://github.com/mbzuai-nlp/ArTST for pluricentric languages like Arabic.
Dialectal Coverage And Generalization in Arabic Speech Recognition
Amirbek Djanibekov1thanks: -These authors contributed equally to this work., Hawau Olamide Toyin [^1] footnotemark: 1 Raghad Alshalan2, Abdullah Alitr2, Hanan Aldarmaki1 1Mohamed bin Zayed University of Artificial Intelligence, Abu Dhabi, UAE 2STC, Riyadh, Saudi Arabia {amirbek.djanibekov,hawau.toyin,hanan.aldarmaki}@mbzuai.ac.ae {rsalshalan,aalatir}@stc.com.sa
1 Introduction
Figure 1: The architecture of SpeechT5/ArTST, which contains an encoder-decoder module and six modal specific pre/post-nets. During self-supervised pre-training (left), quantized tokens are shared across speech and text modalities. Hidden states and latent units are mixed up and used as the inputs of the cross-attention module in the decoder. The fine-tuning stage for ASR is shown on the right. Refer to Ao et al. (2021) for more details.
The advent of large self-supervised acoustic models has transformed speech technology, enabling transfer learning and improving performance for both high-resource and low-resource languages. Prominent examples of such models include various versions of wav2vec Schneider et al. (2019); Baevski et al. (2020), HuBERT Hsu et al. (2021), and SpeechT5 Ao et al. (2021), which have predominantly been trained on English datasets. Their multi-lingual variants, e.g. XLS-R Babu et al. (2021) with 53 and 128 languages, in addition to many models that include both self-supervised and supervised pre-training, such as Whisper Radford et al. (2023) with approximately hundred supported languages, MMS Pratap et al. (2024) with thousands of languages, and UniSpeech Wang et al. (2021), underscore the potential for cross-lingual transfer learning for more inclusive ASR. Yet, while these models indeed show great potential for transfer learning to new languages, even those unseen in training Huang et al. (2013), they remain sub-optimal for specific target languages. A case in point is the Arabic Text and Speech Transformer (ArTST), a model pre-trained exclusively on Arabic, which has demonstrated superior performance for Modern Standard Arabic (MSA), surpassing larger multilingual models like Whisper and MMS in benchmark tests, in addition to establishing a new state-of-the-art (SOTA) performance compared to previous efforts for Arabic ASR. This highlights the advantage of monolingual pre-training when large amounts of unlabeled data for the target language are available. While the model showed some potential for dialectal coverage, it was trained and validated mainly on MSA data, which questions its applicability for spoken dialectal variants of Arabic. Arabic is a pluricentric language Schuppler et al. (2024), diverse in regional variations, and models trained on MSA frequently struggle to adapt to these variations. This limitation is particularly acute given that many Arabic dialects are underrepresented or considered low-resource in speech technology research. Consequently, there is a need for optimized ASR systems that embrace, rather than overlook, the linguistic diversity of the Arabic-speaking world.
In light of these challenges, we conduct various investigations aimed at understanding and enhancing the dialectal diversity and performance of Arabic ASR systems. We focus on three inquiries aimed at optimizing potential strategies for integrating dialectal variation into ASR systems. First, we measure the impact of incorporating a broad collection of Arabic dialects during the model’s pre-training phase. We hypothesize that a wider dialectal foundation could improve the model’s performance across various dialects in the fine-tuning stage. Second, we quantify the comparative effectiveness of dialect-specific fine-tuning versus a more holistic, multi-dialectal fine-tuning strategy. The third question examines the model’s capacity for zero-shot transfer to dialects not explicitly included in fine-tuning. Our key findings from experiments spanning over 12 Arabic dialects are: (1) pre-training with more data and wider dialectal coverage improves performance across most dialectal variants, including MSA, (2) multi-dialectal fine-tuning improves performance for low-resource dialects, but may not be optimal for high-resource dialects, and (3) multi-dialectal pre-training and fine-tuning has higher potential for zero-shot transfer to unseen dialects.
Our pre-training checkpoints and joint models are trained exclusively on open-source data and will be released as open-source models.
2 Related Work
Recent research in Arabic speech recognition, as presented in Hussein et al. (2022), demonstrates the potential of contemporary deep learning techniques in decoding Arabic spoken language. However, this initial success was confined to Modern Standard Arabic (MSA), the formal variant predominant in news broadcasts and official communications. Large-scale multi-lingual ASR models, Whisper Radford et al. (2023) and MMS Pratap et al. (2024), cover many languages within their scope, including Arabic. They utilize language embeddings or adapters to enhance language coverage and performance within the same model, but their performance across languages vary considerably. Toyin et al. (2023) demonstrated state-of-the-art performance in multi-task training for Arabic speech recognition and synthesis, improving over much larger multi-lingual models, but their model was still trained predominantly on MSA. Their findings, however, demonstrate the model’s preliminary potential for processing Arabic dialects. Waheed et al. (2023) explored dialectal fine-tuning of large speech models like Whisper Radford et al. (2023), and XLS-R Babu et al. (2021), demonstrating their effectiveness on Egyptian and Moroccan dialects. The generalizability of these models to other Arabic dialects remains an open question.
3 Research Methodology
Our objective is to optimize state-of-the-art ASR models for dialectal diversity. To that end, we start with the current state-of-the-art model for Arabic, ArTST Toyin et al. (2023), as the foundation model for our investigation. Figure 1 illustrates the high-level architecture of ArTST for self-supervised pre-training and fine-tuning. This model is based on the SpeechT5 approach Ao et al. (2021), and supports multi-modal fine-tuning. The model was pre-trained on the MGB2 Ali et al. (2016) dataset, which consists of newswire data, mainly in MSA, with a small subset of dialectal variants. The model showed some potential for dialectal coverage in ASR, in addition to demonstrating SOTA performance in the task of dialect identification Ali et al. (2019). In this work, we focus on ASR performance across dialects, and attempt to quantify the effect of various design decisions in the development of such ASR models. To understand the factors that enable wider dialectal coverage, we explore the following questions:
  1. 1.
    Is pre-training on dialectal data essential for improving down-stream performance?
  2. 2.
    Is it better to fine-tune ASR models on multiple dialects or fine-tune only on the target dialect?
  3. 3.
    Can we achieve reasonable zero-shot performance on unseen dialects?
  4. 4.
    What is the effect of data augmentation on dialectal fine-tuning performance?
The following sections describe our experimental settings, results, and findings for each of these empirical questions.
Dataset Dialect Hours Words v1 v2
ASC MSA 3.7 hrs 20.58 K
CommVoice mix 133.24 hrs 494.83 K
ClArTTS CA 12 hrs 76.31 K
MGB2 mostly MSA 1000 hrs 7.31 M
QASR mostly MSA 2000 hrs 13.33 M
MGB3[ASR] EGY 2.83 hrs 18.93 K
MGB3[ADI] EGY 11.15 hrs
GLF 8.92 hrs
LAV 9.27 hrs
MSA 9.39 hrs
NOR 9.49 hrs
MGB5[ADI] ALG 115.7hrs
EGY 451.1 hrs
IRA 815.8 hrs
JOR 25.9 hrs
KSA 186.1 hrs
KUW 108.2 hrs
LEB 116.8 hrs
LIB 127.4 hrs
MAU 456.4 hrs
MOR 57.8 hrs
OMA 58.5 hrs
PAL 121.4 hrs
QAT 62.3 hrs
SUD 47.7 hrs
SYR 119.5 hrs
UAE 108.4 hrs
YEM 53.4 hrs
MADAR MOR ALG TUN 532.37K
LIB EGY LEV
IRA GLF YEM
NADI ALG BAH EGY 702.67K
IRA JOR KUW
LEB LIB MOR
OMN PAL QAT
SAU SUD SYR
TUN UAE YEM
Table 1: Summary of Dataset Statistics for Pre-Training: Hours of Audio, Word Counts, and Associated Dialects.
3.1 Terminology
For the rest of the paper, we will refer to Arabic variants using abbreviations. The categories below are based on regions and countries, and do not reflect any official classification of dialectal families:
MSA:
Modern Standard Arabic. This is a common official variant of Arabic used in news, books, and education. CA: Classical Arabic. This is an old variant of Arabic found on religious texts and old books. It resembles MSA, but also contains outdated lexical items and structures.
GLF:
A broad category of dialects spoken in the Arabian Peninsula, in particular the Gulf region, which, in our data sources, include SAU: Saudi, KUW: Kuwait, UAE, OMA: Oman, QAT: Qatar, IRA: Iraq, and YEM: Yemen.
LEV:
Levantine dialects, which, in our data sources, include SYR: Syria, JOR: Jordan, LEB: Lebanon, and PAL: Palestine.
NOR:
North African dialects, including EGY: Egypt, TUN: Tunisia, MOR: Morocco, ALG: Algeria, MAU: Mauritania, and SUD: Sudanese.
3.2 Pre-Training Data & Settings
To examine the effect of pre-training data coverage on downstream performance, we pre-trained ArTST from scratch [^2] We used the scripts and configurations provided in github.com/mbzuai-nlp/ArTST on both MSA and dialectal data. We sourced our data from various dataset: MGB2 Ali et al. (2016), QASR Mubarak et al. (2021) MGB3 Ali et al. (2017), MGB5 Ali et al. (2019), ClArTTS Kulkarni et al. (2023), ASC Halabi et al. (2016), and Common Voice Ardila et al. (2019). We also used MADAR Bouamor et al. (2018) and NADI Abdul-Mageed et al. (2023) text datasets for pre-training. In our experiments, we compare two variants of ArTST:
  • v1: This variant is as described in Toyin et al. (2023), which was pre-trained only on MSA.
  • v2: In this variant, we use a mixture of MSA and dialectal data in pre-training.
See Table 1 for details of the datasets used in pre-training the two variants.
Dataset Dialect Hours Words
QASR mostly MSA 2000 hrs 13.33 M
MGB2 mostly MSA 1000 hrs 7.31 M
MGB3[ASR] EGY 2.83 hrs 18.93 K
MGB5[ASR] MOR 6.74 hrs 56.97 K
SADA Alharbi et al. (2024) SAU 418 hrs 3.25 M
Mixat Al Ali and Aldarmaki (2024) UAE 15 hrs 57.94 K
TARIC-SLU Mdhaffar et al. (2024) TUN 8 hrs 72.00 K
ParallelCorp Almeman et al. (2013) MSA 32 hrs 30.66K
GLF 32 hrs 27.26K
LEV 32 hrs 18.43K
EGY 32 hrs 48.56K
MASC MSA 612.28 hrs 3.80 M
SAU 452.24 hrs 301.92 K
SYR 211.33 hrs 1.06 M
EGY 175.36 hrs 1.03 M
JOR 42.21 hrs 330.83 K
LEB 25.20 hrs 155.76 K
IRA 17.37 hrs 121.12 K
TUN 12.17 hrs 34.34 K
Multiple 10.57 hrs 80.08 K
UAE 9.87 hrs 6.42 K
MOR 8.60 hrs 58.38 K
PAL 6.17 hrs 45.35 K
KUW 4.04 hrs 32.37 K
Table 2: Summary of Dataset Statistics for Fine-Tuning: Hours of Audio, Word Counts, and Associated Dialects. Multiple is mix of several dialects not neccessary from the listed dialects (no information from the source)
3.3 Fine-Tuning Data & Settings
The dataset details used for fine-tuning are shown in table 2. Note that, for MGB5 and MGB3, as the data is based on YouTube videos, many of the originally referenced videos are no longer available, so at the time of our experiments, only 2.5 hours of training were available for MGB3 and 2 hours for MGB5. Furthermore, multi dialectal datasets, such as MASC Al-Fetyani et al. (2021), have unbalanced representation of dialects. The high-resource dialects in our collection include SAU, SYR, EGY, and MSA; each has at least 200 hours of transcribed ASR data. UAE, MOR, JOR, LEB, IRA, and TUN have a medium amount of fine-tuning data between 10 and 50 hours. KUW and PAL are low-resource dialects with less than 10 hours of transcribed data in total. Finally, we left ALG, YEM, and SUD from the MASC dataset for zero-shot testing.
Figure 2 illustrates the distribution of dialectal data in pre-training and fine-tuning stages for a direct comparison. We exclude MSA from the figure as it has disproportionally more data than all dialects.
Figure 2: Distribution of dialectal speech data in pre-training and fine-tuning. MSA is not shown.
3.4 Experimental Settings
For partitioning the data into training, development, and test sets, we adhered to the official splits provided with each dataset. We followed the data preparation and training methodology established in the original ArTST implementation. For comprehensive details regarding the model architecture and data preprocessing, readers are directed to Toyin et al. (2023). The pre-training phase was executed on a cluster of 4 A100 GPUs over a duration of 21 days. We used Adam optimizer with a learning rate of $2\times 10^{-4}$, spanning 335K updates, and a warm-up phase of 64K updates. The maximum speech token length was set at 250K (equivalent to 15.625 seconds). The fine-training phase was executed on a cluster with one A100 GPUs over a duration of 7 days (MGB2, QASR, MASC) and 2 days (MGB3, MGB5, etc.). We used Adam optimizer with a tri stage scheduler for learning rate of $6\times 10^{-5}$.
Normalization
Prior to model training, we implemented the same data normalization steps outlined in Toyin et al. (2023). In addition, we applied post-prediction normalization steps before calculating Word Error Rates (WER), following standard practices in Arabic ASR. All reported results reflect post-normalization performance. The normalization script, sourced from a publicly available GitHub repository333github.com/iamjazzar/arabic-nlp/blob/master/normalization/orthographic_normalization.py, performs orthographic standardization of Alef, Yaa, and Taa characters.
4 Effect of Pre-Training Data
In this section, we examine the effect of pre-training on downstream ASR performance. As described in section 3.2, we use two variants of the model, one pre-trained mainly on MSA (v1), and one trained on additional dialectal data (v2). Note that pre-training does not utilize aligned speech and text; it incorporates un-aligned speech and text data for self-supervised learning.
4.1 Benchmarking MSA Performance
We first report results on benchmark datasets to compare the performance of both models against the state of the art. MGB2 is the main benchmark for MSA speech recognition. We evaluate the performance of both v1 and v2 models fine-tuned on MGB2 on Table 3 compared to existing SOTA models. The results confirm that incorporating dialectal data in pre-training does not negatively affect the performance on MSA.
System WER(%) CER(%)
From Hussein et al. (2022):
HMM-DNN 15.80
E2E, CTC + LM 16.90
E2E, Attention + LM 13.40
E2E, CTC , Attention + LM 12.50
ArTST v1+ LM Toyin et al. (2023) 12.78 6.33
ArTST v2 (ours) 12.49 6.44
ArTST v2 (ours) + LM 12.39 6.51
Table 3: Comparing ArTST performance against models reported in Hussein et al. (2022) and Toyin et al. (2023), which include best performing model previously reports on MGB2.
4.2 Benchmarking Dialectal Performance
Tables 5 and 5 show the performance of the models on the dialectal MGB3 (Egyptian) and MGB5 (Morocooan) benchmarks. Each of these benchmarks contain multiple references as dialectal speech has no standard spelling. We report the average and multi-refrence WER for our model variants, and compare against the best model in each challenge, as well as the state-of-the-art model in each benchmark. Each model is first fine-tuned on MSA, then fine-tuned again on the target MGB train set. We also report the results of the large multilingual models: Whisper and MMS, fine-tuned on the same set. We refer to the data the model is previously adapted on as ‘Adaptation’ data. Starting with MSA data before fine-tuning on the target dialect has previously been established as an effective strategy for dialectal ASR Ali et al. (2017).
System Adaptation Fine-Tuning AV-WER MR-WER
Aalto MGB2 MGB3 37.50 29.30
Whisper ComVoice MGB3 39.04 24.92
Fleurs
Covost2
MMS BibleTrans MGB3 100.04 99.92
NewTestamentRec
v1 MGB2 MGB3 37.08 29.39
v2 MGB2 MGB3 33.20 25.28
Table 4: WER(%) on MGB3 Egyptian ASR. Aalto is the best system in the MGB3 challenge Ali et al. (2017)
System Adaptation Fine-Tuning AV-WER MR-WER
RDI-CU MGB2 MGB5 59.40 37.60
Whisper ComVoice MGB5 164.13 227.34
Fleurs
Covost2
MMS BibleTrans MGB5 111.89 102.30
NewTestamentRec
v1 MGB2 MGB5 49.39 27.95
v2 MGB2 MGB5 48.91 28.02
Table 5: WER(%) on Moroccan ASR. RDI-CU is the best system in the MGB5 challenge Ali et al. (2019)
Dataset Zero-Shot Fine-Tuning
Whisper MMS v1 v2 v1 v2
Mixat 56.01 61.51 39.1 35.05 34.29 33.40
TARIC-SLU 138.14 93.54 107.56 106.46 14.70 14.80
ParallelCorp 99.17 83.16 128.72 141.92 9.57 9.31
SADA 82.16 78.28 39.41 29.77 39.24 29.91
MASC Al-Fetyani et al. (2021)
SAU 48.39 65.30 61.23 58.72 27.40 27.33
SYR 26.65 33.21 21.99 18.37 18.64 17.42
EGY 41.73 66.04 50.87 47.17 38.47 36.43
JOR 28.65 54.63 61.23 34.97 19.72 21.08
LEB 40.95 64.58 35.65 42.66 30.01 28.05
IRA 41.69 59.33 50.46 48.03 31.10 34.64
TUN 47.45 60.58 50.37 46.67 19.26 18.52
MOR 65.87 80.84 78.92 66.87 47.59 49.40
PAL 53.20 83.72 77.94 73.53 55.88 53.53
KUW 36.00 81.71 64.74 52.02 50.29 46.24
Table 6: WER (%) in zero-shot and fine-tuning settings. v1 and v2 models are first fine-tuned on MGB2 (MSA).
In MGB3, dialectal pre-training (v2) results in about 4% absolute reduction in WER. Smaller improvement in terms of Average WER is observed for MGB5, where there is no clear advantage observed using v2. We speculate that this difference is a result of the pre-training having a lot more Egyptian than Moroccan data (see Figure 2).
4.3 Zero-Shot & Fine-Tuning Results
To quantify the effect of dialectal pre-training, we compare fine-tuning performance using the same data sets on both v1 and v2. We first fine-tune models on MSA data using the MGB2 benchmark dataset. We test the model performance on other dialects without dialectal fine-tuning (zero-shot) and with dialectal fine-tuning. The results are shown in Table 6. On average, we see improvement in performance in both zero-shot and fine-tuning experiments using v2 compared to v1. We also see that both v1 and v2 perform better than Whisper and MMS in zero-shot settings in most cases. There are some exceptions, such as in KUW, where Whisper performs better than all other models, including the fine-tuned models, but in all other cases v2 performs better. This underscores the advantage of monolingual models compared to multi-lingual performance, as observed in Toyin et al. (2023) and Radford et al. (2023). In addition, the results underscore the importance of dialectal coverage in pre-training: the cases where v2 does not perform better than v1 are all dialects for which pre-training data are limited, such as TUN (no pre-training data) and JOR (smallest dialect size in pre-training).
We conclude from this section that utilizing more diverse data during the unspervised pre-training stage is recommended for optimizing downstream tasks across more dialectal variants.
5 Joint Models & Dialect ID
We assess the relative effectiveness of individual dialectal fine-tuning compared with joint dialect fine-tuning. To that end, we joined multiple dialectal train sets as shown in Table 7.
From MASC, we excluded ALG, YEM, SUD for zero-shot evaluation. The resulting joint corpus consists of 12 dialects including MSA, with approximately 1,501 training hours in total. We train a single joint model using this data.
Dialect Hours Words Source
MSA 612.28 hrs 3.80 M MASC
SAU 452.24 hrs 301.92 K SADA, MASC
SYR 211.33 hrs 1.06 M MASC
EGY 175.36 hrs 1.03 M MGB3, MASC
JOR 42.21 hrs 330.83 K MASC
LEB 25.20 hrs 155.76 K MASC
IRA 17.37 hrs 121.12 K MASC
TUN 12.17 hrs 34.34 K TARIC-SLU, MASC
UAE 9.87 hrs 6.42 K Mixat, MASC
MOR 8.60 hrs 58.38 K MASC
PAL 6.17 hrs 45.35 K MASC
KUW 4.04 hrs 32.37 K MASC
Table 7: Datasets used to train the joint model.
Fine-Tuning Data zero-shot Fine-Tuning Joint Dialects
[…]
SAU 58.72 43.41 27.33 29.41 30.56 29.94
SYR 18.37 16.20 17.42 19.20 22.41 20.30
EGY 47.17 38.78 36.43 45.17 61.06 46.79
JOR 34.97 25.42 21.08 19.63 21.49 20.11
LEB 42.66 40.51 28.05 28.22 29.43 26.89
IRA 48.03 40.27 36.10 29.33 31.75 30.83
TUN 46.67 45.93 26.67 37.23 28.47 27.74
MOR 66.87 55.42 56.63 57.49 53.89 49.10
PAL 73.53 45.59 53.53 46.22 43.90 44.48
KUW 52.02 45.09 46.24 35.43 39.43 37.71
MSA 21.09 16.78 15.34 11.59 12.66 12.09
Macro Average 46.37 37.58 33.17 32.63 34.09 31.45
Table 8: WER (%) of various models compared with joint dialectal fine-tuning with different dialect ID strategies
5.1 Dialect ID
We trained another model with the aforementioned joint dataset, but with the inclusion of explicit dialect identifiers. We augmented the dictionary with special tokens for dialect IDs, and used them to prepend the decoding string:
<S> DIALECT T1 T2 ... Tn </S>
For inference, we experimented with two strategies: (1) Transcribing with dialect forcing, where we manually add the dialect ID to condition the decoder output; the decoder is forced to start predictions with the tokens <S> DIALECT . (2) Transcribing with dialect inference, where we let the model predict the dialect token automatically. We use this approach for zero-shot ASR on unseen dialects.
The results of the models trained with joint dialects compared to models trained on MGB2 and QASR are shown in Table 8. Note that both MGB2 and QASR contain mostly MSA, but also a small amounts of various dialects, but their exact distribution is unknown. We also show the fine-tuning results from Table 6 for easy comparison. We see that joint modeling results in improvement for low-resource dialects, including: JOR, TUN, and KUW, but degrades performance of the high-resource SYR and EGY dialects. Interestingly, dialect forcing, results in average performance similar to target fine-tuning, and dialect inference resulted in the best performance overall, followed by the model without dialect ID. We surmise that the model learns dialectal patterns that do not perfectly align with the dialect ID as indicated in the training data. Since the dialect IDs are coarse country-level approximations, letting the model infer the dialect that best aligns with the speech is the best approach for most cases. Many dialectal sets, such as SYR and SAU, contain a lot of MSA utterances that are incorrectly identified as dialectal, and low-resource dialects, such as KUW, are predicted as their closest high-resource variant, such as SAU.
Figure 3 illustrate dialect inference errors. Note that the number of errors is proportional to the test data size. The overall dialect identification performance is around 90%.
[…]
6 Effect of Data Adaptation
Fine-Tuning Adaptation MGB2 QASR
MGB3 136.52 25.28 19.53
MGB5 94.84 28.02 27.58
TARIC-SLU 30.46 14.80 14.48
ParallelCorp 28.97 9.31 9.08
SADA 29.77 29.91 29.55
Mixat 100.0 33.40 35.21
Table 9: WER (%) of fine-tuned models on various datasets with different adaptation methods: —: no adaptation, MGB2, or QASR.
dialects Fine-Tuning Set
SAU(%) SYR(%) EGY(%) JOR(%) LEB(%) IRA(%) TUN(%) MOR(%) PAL(%) KUW(%) UAE(%)
test set SAU 27.33 53.88 63.24 43.95 56.64 46.10 49.36 50.02 45.48 46.85 46.14
SYR 19.54 17.42 23.91 16.88 25.71 16.85 24.36 17.58 16.86 17.38 17.53
EGY 38.07 53.43 36.43 40.08 51.61 41.37 45.58 43.88 40.41 41.88 40.38
JOR 22.88 30.37 34.14 21.08 28.11 28.68 32.96 30.50 25.25 27.88 28.05
LEB 41.03 42.53 53.23 39.07 28.05 41.23 48.53 42.34 41.55 42.14 40.97
IRA 35.18 49.10 56.91 40.47 46.72 36.10 47.84 42.55 40.76 41.00 42.16
TUN 44.44 57.78 51.85 46.67 45.19 45.19 26.67 44.44 47.41 46.67 47.41
MOR 59.04 71.69 74.70 54.82 69.88 57.23 74.70 56.63 56.02 59.64 57.83
PAL 48.53 66.18 62.65 52.35 60.59 58.53 60.00 63.24 53.53 58.53 60.59
KUW 26.59 59.54 78.03 43.93 67.63 46.82 53.18 51.45 44.51 46.24 50.87
held-out ALG 50.21 60.09 73.61 52.58 62.02 57.51 57.51 59.01 52.79 58.37 57.30
SUD 40.89 64.97 64.32 53.15 64.86 54.12 56.51 58.79 52.39 56.18 55.53
YEM 38.40 42.16 38.49 34.68 39.69 34.92 30.68 35.73 33.06 34.68 34.11
[…]
ALG SUD YEM
WER (%) WER (%) WER (%)
[…]
v2 Joint 45.20 40.69 33.08
v2 joint (w. dialect inference) 47.12 40.15 31.84
Table 11: WER of various models on held-out dialects.
In the above experiments, we followed the strategy of initializing the models by first fine-tuning on MSA data. In most cases, we used MGB2 as the base model, following previously established results on Egyptian ASR Ali et al. (2017). This adaptation approach is meant to enhance the performance on low-resource dialects, facilitating faster convergence with limited training samples. However, as pre-training on more diverse sets proved to be effective, adaptation on more diverse data is also likely to be fruitful. As observed in Table 8, models trained on QASR resulted in far better zero-shot performance, approaching the performance of joint-dialects models. This is attributed to the fact that QASR is both larger in size and known to have more dialectal data compared to MGB2 (but both have no documented statistics of dialectal coverage). To validate this observation, we experimented with dialectal fine-tuning adapted from two variants: one based on MGB2 and one based on QASR Mubarak et al. (2021), followed by dialect-specific fine-tuning. Table 5 shows fine-tuning results with no adaptation (directly fine-tuning on the target dialect), compared with starting from MGB2 or QASR. First, our results corroborate the previous findings that adapting models from MSA results in large reduction in error rates. In all except the Mixat dataset, starting from QASR results in better performance compared to MGB2. However, the difference is negligible except on MGB3 Egyptian set (around 10% absolute WER reduction). There are two factors that we speculate underline this result: The small size of the MGB3 set, and the existence of Egyptian dialect in the QASR corpus more substantially than the other dialects. Overall, using the QASR dataset as a basis for adapting dialectal models is recommended as it improves or maintains performance.
7 Cross-Dialectal Performance
Table 10 shows the cross-dialectal performance, where models trained on a single target dialect are tested on other dialects, including the three held-out sets: ALG, SUD, and YEM. We also show the performance of zero-shot and joint models on the held-out dialects in Table 11. In most cases, the best performance is achieved by the model trained on the same target dialect (the diagonal in Table 10). However, for low-resource dialects, like KUW and PAL, the model trained on SAU achieved the lowest WER. This is likely a result of the large size of the SAU train set and the wide geographical area and dialectal variants it covers. The joint models perform best on the held-out dialects. Overall, the performance on the held-out sets is on a par with the low-resource dialects, with WER above 30%. Curiously, all models perform well on the SYR test set; upon close inspection, we found that the set consists mostly of MSA utterances, which explains the result since all models are adapted from MSA.
8 Conclusions
We presented the largest study on dialectal Arabic ASR to empirically demonstrate the effect of various training paradigms on ASR performance. We compared models pre-trained with and without dialects, in high, low, and medium-resource settings, in addition to zero-shot. We find that overall, dialectal pre-training improves performance in zero-shot and low-resource cases, and mostly maintains performance on MSA and high-resource dialects. We also find that all dialects benefit from adaptation of models pre-fine-tuned on MSA, and this effect is most noticeable for low and medium-resource dialects. Adaptation from larger and more varied datasets, like QASR, results in improved performance, particularly in ’zero-shot’ settings. The latter is likely influenced by the presence of dialectal data in QASR, so it is not pure zero-shot as the exact composition of the dataset is not documented. We experimented with more controlled multi-dialectal fine-tuning, where we joined the train sets of 12 dialects. In this case, we observe additional improvements compared to QASR in spite of having smaller size overall, where we achieve on average at least the same performance as the target-dialect fine-tuning setting, and the best performance on held-out dialects. Interestingly, while using dialect ID in decoding is effective, forcing the dialect ID results in worse performance compared to dialect inference. Finally, while joint training results in improved performance for the medium and low-resource dialects, target-dialect fine-tuning is more effective for high-resource dialects. In light of the above experiments and results, we recommend the following: (1) include as many dialectal variants as possible in the pre-training stage, (2) fine-tune ASR on large sets of joint dialects, preferably with the dialect ID inference, then (3) fine-tune separate models for high-resource dialects with more than 100 hours of labeled data. To enable easier adoption and further experiments, we will release the pre-trained v2 checkpoint, the fine-tuned MGB2 and QASR models, and the joint model with dialect tokens.
Limitations
One of the limitations in dialect-related work is the coarse classification of dialect IDs; dialects in our datasets are classified by regions or countries, whereas actual dialectal variations are far more fine-grained. For example, the Saudi dataset, SADA, covers a large geographical area and many dialects, but it is considered as one dialect based on our classification. Moreover, the way the datasets are collected do not guarantee that the data are indeed dialectal. For instance, with manual inspection of the Syrian test and dev sets from MASC, we observed that all instances are in MSA rather than Syrian dialects. In addition, Arabic dialects are spoken varieties that do not have a standard spelling system. This results in large variations in transcriptions, but standard WER does not account for these variations, resulting in more pessimistic results. With the exception of the MGB3 and MGB5 benchmarks where we report average and multi-reference WER across 4 references, all datasets have only a single reference. Code-switching is also observed in some datasets, such as Mixat, but the model has limited ability to recognize multi-lingual code-switching as it is only trained on Arabic. Given small amount of English code-switching in the pre-training data, the model does have limited code-switching recognition, but it can be improved by intentionally incorporating code-switching datasets for pre-training and fine-tuning. The combination of multi-lingual and multi-dialectal training has not been explored in this work.
Appendix A Inference examples
Figure 4: Examples of dialectal recognition after targeted fine-tuning, following MGB2 adaptation.
Figure 5: Examples from the Mixat & TARIC-SLU dataset illustrating the zero-shot performance of v1 and v2 fine-tuned on MGB2 vs. v2 fine-tuned on QASR.
Figure 6: Example of ASR outputs on held-out dialects from dialect-specific fine-tuning.
Table 4 lists examples of ASR outputs using the dialect-specific fine-tuned models. Note that the ‘errors’ in SAU, EGY, and JOR examples are in fact alternative spellings. Figure 5 compares zero-shot v1 and v2 (adapted from MGB2) and v2 (adapted from QASR) performance from the Mixat (UAE) and TARIC (Tunisia) datasets.